OpenStudy (anonymous):

Please help! How many solutions does the equation have? A. one solution B. infinite solutions C. no solution

OpenStudy (anonymous):

What equation?

OpenStudy (anonymous):

equation?

OpenStudy (anonymous):

oops, guess it didn't copy. 4(2x+3)=2(3x-4)

OpenStudy (leader):

Distribute 8x+12=6x-8 2x=-20 x=-10

OpenStudy (anonymous):

Thank you! @Leader

OpenStudy (leader):

No problem :)